Translate the following into a quadratic equation, and solve it: The length of a rectangular garden is 3 times it's width; if the area of the garden is 75 square meters, what are it's dimensions?: Translate the following into a quadratic equation, and solve it: The length of a rectangular garden is 3 times it's width; if the area of the garden is 75 square meters, what are it's dimensions?
Answer by checkley77(3394)

L*W=AREA
3X*X=75
3X^2=75
X^2=75/3
X^2=25
X=SQRT25
X=5 ANSWER FOR THE WIDTH.
3*5=15 ANSWER FOR THE LENGTH.
PROOF:
5*15=75
75=75